OpenStudy (anonymous):

what is the limit of (2-(e)^x/2)^4/x as x approaches 0??

OpenStudy (anonymous):

is it infinite?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \ \ (2-e^{\frac{x}{2}})^{\frac{4}{x}}\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

letting x=0 it becomes \[1^\infty\]undefined

OpenStudy (anonymous):

?

OpenStudy (anonymous):

if I apply the LHR?

OpenStudy (anonymous):

yeah LHR so let\[y= (2-e^{\frac{x}{2}})^{\frac{4}{x}}\]\[\ln y=\frac{4}{x} \ln(2-e^{\frac{x}{2}})=4\frac{\ln(2-e^{\frac{x}{2}})}{x}\]apply the LHR for right hand side

OpenStudy (anonymous):

that would be?

OpenStudy (anonymous):

can u apply LHR for\[\frac{\ln(2-e^{\frac{x}{2}})}{x}\]as x--->0

OpenStudy (anonymous):

\[\frac{ \frac{ x }{ 2 }(-e ^{\frac{ x }{ 2}}) }{ 2-e ^{\frac{ x }{ 2 }}}\]?

OpenStudy (shubhamsrg):

well for 1^(infinity) forms,,remember this for convenience .. ans would be e^(f(x).g(x)) where we had something like this : (1+f(x))^g(x) ,,here f(x)-->0 and g(x)-->infinity..

OpenStudy (shubhamsrg):

in your case, f(x) = 1- e^(x/2) and g(x) = 4/x ans would be: e^[ (x-->0) (1-e^(x/2) * 4/x ] you may apply LH rule here,,

OpenStudy (shubhamsrg):

was i able to explain well ?

OpenStudy (anonymous):

still don't get it